# Data-Locality-Aware Task Assignment and Scheduling for Distributed Job Executions

### v1

@@ II Problem Formulation
We denote by $\mu_{m}^{c}$ the profiled number of tasks for job $c$ that can be processed at server $m$ within a time slot. We define the busy time of each server $m$ just before job $c$’s arrival, denoted by $b_{m}^{c}$, as the number of time slots required to finish all the outstanding job tasks in the corresponding queue. The busy time $b_{m}^{c}$ can be estimated using the following equation:
@@
Our task is to make assignment and scheduling decisions for each arriving job. The objective is to minimize the completion time of each job. For each job, the completion time is defined as the time between its arrival and the completion of its last processed task. This objective is equivalent to minimizing the completion time of the last processed task issued by the job. We study the problem in two scenarios:
  • FIFO queues. The outstanding job tasks are queued in servers in a FIFO manner. The key is to assign the tasks of each arriving job to their available servers in a way that minimizes the job completion time by keeping the servers’ estimated busy time as balanced as possible.
@@ III-A1 Programming in a Bipartite Graph
Recall that $b_{m}^{c}$ is the estimated busy time of server $m$ before the assignment of tasks in $\mathcal{T}_{c}$, calculated by (2). In $\mathcal{P}$, $n_{m}^{k}$ is the number of time slots required to process the tasks in group $\mathcal{T}_{c}^{k}$ at server $m$, given the profiled processing capacity $\mu_{m}^{c}$. Each $n_{m}^{k}$ is associated with an edge $(k,m)$ in the bipartite graph. $\Phi_{c}$ is the estimated completion time of job $c$. The first constraint in (6) means that any available server whose estimated busy time is larger than $\Phi_{c}$ should not participate in the assignment of tasks in $\mathcal{T}_{c}$. This constraint ensures balance among the participating servers. The second constraint in (6) ensures that all the tasks in $\mathcal{T}_{c}$ can be processed. The program $\mathcal{P}$ can be solved optimally by commercial solvers, e.g., CPLEX. [^1] https://www.ibm.com/products/ilog-cplex-optimization-studio
@@ III-A3 The Matching-based Algorithm
By introducing $\Phi_{c}^{-}$ and $\Phi_{c}^{+}$, the search space of $\Phi_{c}$ is greatly narrowed. However, $\mathcal{P}$ is still non-linear since the first constraint in (6) is piecewise. To solve $\mathcal{P}$ efficiently, we divide $\left[\Phi_{c}^{-},\Phi_{c}^{+}\right]$ into several disjoint sub-intervals based on the estimated busy times of available servers. The division is illustrated in Fig. 1. Specifically, we sort the estimated busy time $b_{m}^{c}$ of each server $m\in\bigcup_{k\in\mathcal{K}_{c}}\mathcal{S}_{c}^{k}$ in ascending order, and we denote the sorted values by $b_{1}^{\prime},b_{2}^{\prime},\ldots$. We then denote by $i$ ($j$) the minimal (maximal) index of these values such that $b_{i}^{\prime}\geq\Phi_{c}^{-}$ ($b_{j}^{\prime}\leq\Phi_{c}^{+}$). Then, we can solve $\mathcal{P}$ with the search space of $\Phi_{c}$ falling into each sub-interval $[\Phi_{c}^{-},b_{i}^{\prime}),[b_{i}^{\prime},b_{i+1}^{\prime}),\ldots,[b_{j}^{\prime},\Phi_{c}^{+}]$ in turn. Each of these problems is transformed into a linear integer program since the first constraint of (6) is no longer piecewise. The sub-intervals are checked one by one. If $\mathcal{P}$ is solvable within the current sub-interval, then its solution is the optimal one since the remaining sub-intervals cannot contain a smaller $\Phi_{c}$.
@@ 3             Estimate $b_{m}^{c}$ by (2)
6             Sort the servers in $\mathcal{S}_{c}^{k}$ in ascending order of their estimated busy times, i.e., $b_{m}^{c}$
7             for each server $m$ in the sorted $\mathcal{S}_{c}^{k}$ do
8                   if server $m$ is not the last one in $\mathcal{S}_{c}^{k}$ then
@@ III-B1 Algorithm Design
for each $m\in\mathcal{S}_{c}^{k}$. WF repeats this procedure for each task group sequentially, performing like filling water, ensuring that the participating servers have the same estimated busy time. Fig. 2 illustrates the idea of water-filling
Figure 2: Assign tasks in $\mathcal{T}_{c}^{k}$ to their available servers in a water-filling manner. In (i), only a subset of $\mathcal{S}_{c}^{k}$, includes servers $1,...,m^{*}$, participate in the assignment of $\mathcal{T}_{c}^{k}$; In (ii), every server in $\mathcal{S}_{c}^{k}$ participates in the assignment of $\mathcal{T}_{c}^{k}$.
@@ III-B2 Approximation Analysis
where $I:=I\big{(}c,\{b_{m}^{c}\}_{m}\big{)}$ is an arrival instance of a new job $c$ and the estimated busy time of each server at job $c$’s arrival. $\text{OPT}(I)$ and $\text{WF}(I)$ are respectively the maximum estimated busy times of participating servers after the tasks of job $c$ are assigned by the optimal algorithm and WF.
@@
holds for each $k\in\mathcal{K}_{c}$, since $\mathcal{T}_{c}^{k}\subseteq\mathcal{T}_{c}$.
@@
Figure 6: Visualization of Case I. In (i), $\max_{m\in\Omega_{c}^{k+1}}b_{m}^{c}(k+1)>\textsc{WF}_{k}$. In (ii), $\max_{m\in\Omega_{c}^{k+1}}b_{m}^{c}(k+1)\leq\textsc{WF}_{k}$.
@@
With this result, we prove $\textsc{WF}_{k+1}-\textsc{WF}_{k}\leq\textsc{OPT}$ by comparing $\Omega_{c}^{k+1}$ and $\beth_{c}^{k+1}$, where $\beth_{c}^{k+1}\subseteq\mathcal{S}_{c}^{k+1}$ is the set of participating servers for the assignment of $\mathcal{T}_{c}^{k+1}$ by $\textsc{OPT}_{k}$.
@@
In $\textsc{OPT}_{k+1}$, $\textsc{OPT}_{k+1}$ time slots of all the participating servers $\beth_{c}^{k+1}$ can accommodate all the tasks of $\mathcal{T}_{c}^{k+1}$ and the outstanding tasks of previous jobs queuing at the servers in $\beth_{c}^{k+1}\cap\Omega_{c,>}^{k+1}$. Thus,
@@ III-C1 Algorithm Design
After initialization, RD enters the deletion phase. To minimize the completion time of job $c$, RD removes redundant task replicas while keeping the estimated busy time of participating servers as balanced as possible. In each deletion iteration, RD first identifies the participating server(s) with the largest estimated busy time, referred to as the target server(s), denoted by $m^{\star}$. For example, in the first deletion iteration (Fig. 8(i)), the target server is server 2. RD then removes up to $\mu_{m^{\star}}^{c}$ task replicas from the target server to reduce its estimated busy time by one time slot. In each deletion iteration, RD removes task replicas with the largest numbers of copies.
@@
RD exits the deletion phase when all the tasks in the target server(s) have only one replica, indicating that the completion time of job $c$ cannot be reduced further. In Fig. 8(viii), the deletion phase terminates since all the tasks in target servers 1 and 5 have no other replicas. In the final phase, RD removes redundant task replicas from the remaining participating servers so that each task is processed by only one server. This removal procedure operates similarly to the deletion phase, identifying and removing replicas from the server with the largest estimated busy time among the remaining servers. For example, in Fig. 8(ix), RD deletes the replicas of the yellow and green tasks from server 4. This is done to keep the estimated busy time of all the servers as small and balanced as possible.
RD assigns tasks to servers from a global perspective because it scans all available servers when selecting a task replica to delete. In contrast, WF minimizes the completion time of the arrived job locally by partitioning tasks into task groups and balancing only the available servers within each task group in each iteration. Intuitively, RD can outperform WF in most cases, but RD’s procedure is more complex to execute.
@@ III-C2 Complexity Analysis
In the implementation of RD, we use a priority queue to maintain all servers, prioritizing by their estimated busy times (and break ties by their initial estimated busy times). Additionally, for each server $m$, we use a priority queue to maintain the assigned tasks, prioritizing by their numbers of replicas. In the worst case, each task is replicated on all servers, requiring $\mathcal{O}(|\mathcal{T}_{c}|\cdot M)$ deletions. Moreover, each deletion affects the priority queues of the servers containing replicas of the deleted task. Thus, each deletion leads to updating the priority queue of up to $M$ servers, resulting in a complexity of $\mathcal{O}(M\cdot\log|\mathcal{T}_{c}|)$. Therefore, the total time complexity of RD is $\mathcal{O}(M^{2}\cdot|\mathcal{T}_{c}|\cdot\log|\mathcal{T}_{c}|)$. We experimentally compare the performance of RD with WF and other algorithms in Sec. V, and leave its theoretical analysis as future work.
@@ IV Algorithms with Job Reordering
In this section, we consider an extended scenario in which the outstanding job tasks queued on each server can be prioritized and reordered. By adjusting the execution order of outstanding jobs and reassigning their remaining tasks, their average completion time can be further reduced.
@@
Assume that there are $p$ sorted outstanding jobs in $\mathcal{Q}_{c}$. For each job $c^{\prime}\in\mathcal{O}_{c}\backslash\mathcal{Q}_{c}$, we estimate the remaining time to complete job $c^{\prime}$ as $\Phi_{c^{\prime}}$, assuming that job $c^{\prime}$ is the $(p+1)$-th job in the new order. Here, $\Phi_{c^{\prime}}$ is derived by WF with the estimated busy times of servers, given the assignment of $p$ jobs in $\mathcal{Q}_{c}$ and the number of unprocessed tasks of job $c^{\prime}$. We denote by $l$ the index of the job with the minimal $\Phi_{l}$ among all the outstanding jobs explored. When exploring each job $c^{\prime}\in\mathcal{O}_{c}\backslash\mathcal{Q}_{c}$, we first compute a lower bound $\Phi_{c^{\prime}}^{-}$ on the completion time of job $c^{\prime}$, as indicated by (8) and (9) in Section III-A2. If $\Phi_{c^{\prime}}^{-}\geq\Phi_{l}$, the exploration of job $c^{\prime}$ terminates. Otherwise, we calculate $\Phi_{c^{\prime}}$ by WF and update the best-explored job $l$ if $\Phi_{c^{\prime}}<\Phi_{l}$. The reason is that if $\Phi_{c^{\prime}}^{-}>\Phi_{l}$, letting job $c^{\prime}$ be the $(p+1)$-th job can never lead to a lower average job completion time. This technique, which significantly reduces the computation overhead on finding the $(p+1)$-th job in the new order, is named early-exit.
@@ V-A Experimental Setup
Job Traces. We use the Alibaba cluster trace program [8] to drive the simulation. We extract a segment from the dataset batch_task.csv in cluster-trace-v2017 that contains 250 jobs. There are a total of 113,653 tasks in these jobs. We derive job arrivals and task durations from the timestamps of the recorded task events. We scale the inter-arrival times of the jobs to simulate different levels of system utilization, ranging from 25% to 75%. The default number of servers is 100. For each server, its computing capacity for each job is randomly generated between 3 and 5 by default.
@@ V-B Experimental Results
Compared with NLIP, OBTA achieves a significant improvement in efficiency. From Figs. 10-12, we observe that OBTA and NLIP have fairly close performance of job completion time since both are theoretically optimal for balanced task assignment. By narrowing the search space and dividing it into subranges, OBTA reduces the computation overhead by nearly half compared to NLIP. This verifies our contribution in Section III-B.
@@
Figure 14: Average job completion time, and CDF of job completion times when $\alpha=2$ under 75% system utilization for different computing capacities.
@@
Figs. 13-14 show the job completion times for varying numbers of available servers and computing capacities, respectively. These experiments are conducted with $\alpha=2$ and a system utilization of 75%, as we are particularly interested in the performance when resource contention is high. From Fig. 13, we observe that a larger number of available servers decreases job completion times. More available servers provide greater flexibility in task assignment, allowing tasks to be distributed in a manner that better balances the load among servers and reduces job completion times. Fig. 14 demonstrates that increasing computing capacity also reduces job completion times. This is expected, as higher computing capacities enable more tasks to run in parallel, thereby decreasing job completion times. Notably, the relative performance of the algorithms remains largely unchanged across different numbers of available servers and computing capacities.
@@ Acknowledgments
This work was supported in part by the Singapore Ministry of Education under Academic Research Fund Tier 2 Award MOE-T2EP20121-0005, the National Science Foundation of China under Grants 62125206 and U20A20173, the National Key Research and Development Program of China under Grant 2022YFB4500100, and the Key Research Project of Zhejiang Province under Grant 2022C01145.

### v3

@@ II Problem Formulation
We denote by $\mu_{m}^{c}$ the profiled number of tasks for job $c$ that can be processed at server $m$ within a time slot. We define the initial busy time of each server $m$ just before job $c$’s arrival, denoted by $b_{m}^{c}$, as the number of time slots required to finish all the outstanding job tasks in the corresponding queue. The initial busy time $b_{m}^{c}$ can be estimated using the following equation:
@@
Our task is to make assignment and scheduling decisions for each arriving job. The objective is to minimize the completion time of each job. For each job, the completion time is defined as the time duration between its arrival and the completion of its last processed task. This objective is equivalent to minimizing the completion time of the last processed task issued by the job. We study the problem in two scenarios:
  • FIFO queues. The outstanding job tasks are queued in servers in a FIFO manner. The key is to assign the tasks of each arriving job to their available servers in a way that minimizes the job completion time by keeping the estimated busy times of the servers as balanced as possible.
@@ III-A1 Programming in a Bipartite Graph
Recall that $b_{m}^{c}$ is the initial estimated busy time of server $m$ before the assignment of tasks in $\mathcal{T}_{c}$, calculated by (2). In $\mathcal{P}$, $n_{m}^{k}$ is the number of time slots required to process the tasks in group $\mathcal{T}_{c}^{k}$ at server $m$, given the profiled processing capacity $\mu_{m}^{c}$. Each $n_{m}^{k}$ is associated with an edge $(k,m)$ in the bipartite graph. $\Phi_{c}$ is the estimated completion time of job $c$. The first constraint in (6) means that any available server whose initial estimated busy time is larger than $\Phi_{c}$ should not participate in the assignment of tasks in $\mathcal{T}_{c}$. This constraint ensures balance among the participating servers. The second constraint in (6) ensures that all the tasks in $\mathcal{T}_{c}$ can be processed. The program $\mathcal{P}$ can be solved optimally by commercial solvers, e.g., CPLEX. [^1] https://www.ibm.com/products/ilog-cplex-optimization-studio
@@ III-A3 The Matching-based Algorithm
By introducing $\Phi_{c}^{-}$ and $\Phi_{c}^{+}$, the search space of $\Phi_{c}$ is greatly narrowed. However, $\mathcal{P}$ is still non-linear since the first constraint in (6) is piecewise. To solve $\mathcal{P}$ efficiently, we divide $\left[\Phi_{c}^{-},\Phi_{c}^{+}\right]$ into several disjoint sub-intervals based on the initial estimated busy times of available servers. The division is illustrated in Fig. 1. Specifically, we sort the initial estimated busy time $b_{m}^{c}$ of each server $m\in\bigcup_{k\in\mathcal{K}_{c}}\mathcal{S}_{c}^{k}$ in ascending order, and we denote the sorted values by $b_{1}^{\prime},b_{2}^{\prime},\ldots$. We then denote by $i$ ($j$) the minimal (maximal) index of these values such that $b_{i}^{\prime}\geq\Phi_{c}^{-}$ ($b_{j}^{\prime}\leq\Phi_{c}^{+}$). Then, we can solve $\mathcal{P}$ with the search space of $\Phi_{c}$ falling into each sub-interval $[\Phi_{c}^{-},b_{i}^{\prime}),[b_{i}^{\prime},b_{i+1}^{\prime}),\ldots,[b_{j}^{\prime},\Phi_{c}^{+}]$ in turn. Each of these problems is transformed into a linear integer program since the first constraint of (6) is no longer piecewise. The sub-intervals are checked one by one. If $\mathcal{P}$ is solvable within the current sub-interval, then its solution is the optimal one since the remaining sub-intervals cannot contain a smaller $\Phi_{c}$.
@@ 3             Estimate $b_{m}^{c}$ by (2)
6             for each server $m\in\mathcal{S}_{c}^{k}$ do
7                   if server $m$ is not the last server in $\mathcal{S}_{c}^{k}$ then
@@ III-B1 Algorithm Design
for each $m\in\mathcal{S}_{c}^{k}$. WF repeats this procedure for each task group sequentially, performing like filling water, ensuring that the participating servers have the same estimated busy time. Fig. 2 illustrates the idea of water-filling.
Figure 2: Assign tasks in $\mathcal{T}_{c}^{k}$ to their available servers in a water-filling manner. In (i), only a subset of $\mathcal{S}_{c}^{k}$, including servers $1,...,m^{*}$, participate in the assignment of $\mathcal{T}_{c}^{k}$; In (ii), every server in $\mathcal{S}_{c}^{k}$ participates in the assignment of $\mathcal{T}_{c}^{k}$.
@@ III-B2 Approximation Analysis
where $I:=I\big{(}c,\{b_{m}^{c}\}_{m}\big{)}$ is an arrival instance of a new job $c$ and the initial estimated busy time of each server just before job $c$’s arrival. $\text{OPT}(I)$ and $\text{WF}(I)$ are respectively the maximum estimated busy times of participating servers after the tasks of job $c$ are assigned by the optimal algorithm and WF.
@@
holds for each $k\in\mathcal{K}_{c}$, since $\mathcal{T}_{c}^{k+1}\subseteq\mathcal{T}_{c}$.
@@
Figure 6: Visualization of Case I. In (i), $\max_{m\in\Omega_{c}^{k+1}}b_{m}^{c}(k+1)\leq\textsc{WF}_{k}$. In (ii), $\max_{m\in\Omega_{c}^{k+1}}b_{m}^{c}(k+1)>\textsc{WF}_{k}$.
@@
With this result, we prove $\textsc{WF}_{k+1}-\textsc{WF}_{k}\leq\textsc{OPT}$ by comparing $\Omega_{c}^{k+1}$ and $\beth_{c}^{k+1}$, where $\beth_{c}^{k+1}\subseteq\mathcal{S}_{c}^{k+1}$ is the set of participating servers for the assignment of $\mathcal{T}_{c}^{k+1}$ by $\textsc{OPT}_{k+1}$.
@@
In $\textsc{OPT}_{k+1}$, $\textsc{OPT}_{k+1}$ time slots of all the participating servers $\beth_{c}^{k+1}$ can accommodate all the tasks of $\mathcal{T}_{c}^{k+1}$ and the outstanding tasks of previous jobs queued at the servers in $\beth_{c}^{k+1}\cap\Omega_{c,>}^{k+1}$. Thus,
@@ III-C1 Algorithm Design
After initialization, RD enters the deletion phase. To minimize the completion time of job $c$, RD removes redundant task replicas while keeping the estimated busy times of participating servers as balanced as possible. In each deletion iteration, RD first identifies the participating server(s) with the largest estimated busy time, referred to as the target server(s), denoted by $m^{\star}$. For example, in the first deletion iteration (Fig. 8(i)), the target server is server 2. RD then removes up to $\mu_{m^{\star}}^{c}$ task replicas from the target server to reduce its estimated busy time by one time slot. In each deletion iteration, RD removes task replicas with the largest numbers of copies.
@@
RD exits the deletion phase when all the tasks in the target server(s) have only one replica, indicating that the completion time of job $c$ cannot be reduced further. In Fig. 8(viii), the deletion phase terminates since all the tasks in target servers 1 and 5 have no other replicas. In the final phase, RD removes redundant task replicas from the remaining participating servers so that each task is processed by only one server. This removal procedure operates similarly to the deletion phase, identifying and removing replicas from the server with the largest estimated busy time among the remaining servers. For example, in Fig. 8(ix), RD deletes the replicas of the yellow and green tasks from server 4. This is done to keep the estimated busy times of all the servers as small and balanced as possible.
RD assigns tasks to servers from a global perspective because it scans all available servers when selecting a task replica to delete. In contrast, WF minimizes the completion time of the arrived job locally by dividing tasks into task groups and balancing only the available servers within each task group in each iteration. Intuitively, RD can outperform WF in most cases, but RD’s procedure is more complex to execute.
@@ III-C2 Complexity Analysis
In the implementation of RD, we use a priority queue to maintain all servers, prioritizing by their estimated busy times (and break ties by their initial estimated busy times). Additionally, for each server $m$, we use a priority queue to maintain the assigned tasks, prioritizing by their numbers of replicas. In the worst case, each task is replicated on all servers, requiring $\mathcal{O}(|\mathcal{T}_{c}|\cdot M)$ deletions. Moreover, each deletion affects the priority queues of the servers containing the replicas of the deleted task. Thus, each deletion leads to updating the priority queue of up to $M$ servers, resulting in a complexity of $\mathcal{O}(M\cdot\log|\mathcal{T}_{c}|)$. Therefore, the total time complexity of RD is $\mathcal{O}(M^{2}\cdot|\mathcal{T}_{c}|\cdot\log|\mathcal{T}_{c}|)$. We experimentally compare the performance of RD with WF and other algorithms in Sec. V, and leave its theoretical analysis as future work.
@@ IV Algorithms with Job Reordering
In this section, we consider an extended scenario in which the outstanding job tasks queued in each server can be prioritized and reordered. By adjusting the execution order of outstanding jobs and reassigning their remaining tasks, their average completion time can be further reduced.
@@
Assume that there are $p$ sorted outstanding jobs in $\mathcal{Q}_{c}$. For each job $c^{\prime}\in\mathcal{O}_{c}\backslash\mathcal{Q}_{c}$, we estimate the remaining time to complete job $c^{\prime}$ as $\Phi_{c^{\prime}}$, assuming that job $c^{\prime}$ is the $(p+1)$-th job in the new order. Here, $\Phi_{c^{\prime}}$ is derived by WF with the estimated busy times of servers, given the assignment of $p$ jobs in $\mathcal{Q}_{c}$ and the number of unprocessed tasks of job $c^{\prime}$. We denote by $l$ the index of the job with the minimal $\Phi_{l}$ among all the outstanding jobs explored. When exploring each job $c^{\prime}\in\mathcal{O}_{c}\backslash\mathcal{Q}_{c}$, we first compute a lower bound $\Phi_{c^{\prime}}^{-}$ on the completion time of job $c^{\prime}$, as indicated by (8) and (9) in Sec. III-A2. If $\Phi_{c^{\prime}}^{-}\geq\Phi_{l}$, the exploration of job $c^{\prime}$ terminates. Otherwise, we calculate $\Phi_{c^{\prime}}$ by WF and update the best-explored job $l$ if $\Phi_{c^{\prime}}<\Phi_{l}$. The reason is that if $\Phi_{c^{\prime}}^{-}>\Phi_{l}$, letting job $c^{\prime}$ be the $(p+1)$-th job can never lead to a lower average job completion time. This technique, which significantly reduces the computation overhead on finding the $(p+1)$-th job in the new order, is named early-exit.
@@ V-A Experimental Setup
Job Traces. We use the Alibaba cluster trace program [8] to drive the simulation. We extract a segment from the dataset batch_task.csv in cluster-trace-v2017 that contains 250 jobs. There are a total of 113,653 tasks in these jobs. We derive job arrivals and task durations from the timestamps of the recorded task events. We scale the inter-arrival times of the jobs to simulate different levels of system utilization, ranging from 25% to 75%. The default number of servers is 100. For each server, its profiled processing capacity for each job is randomly generated between 3 and 5 by default.
@@ V-B Experimental Results
Compared with NLIP, OBTA achieves a significant improvement in efficiency. From Figs. 10-12, we observe that OBTA and NLIP have fairly close performance of job completion time since both are theoretically optimal for balanced task assignment. By narrowing the search space and dividing it into subranges, OBTA reduces the computation overhead by nearly half compared to NLIP. This verifies our contribution in Sec. III-B.
@@
Figure 14: Average job completion time, and CDF of job completion times when $\alpha=2$ under 75% system utilization for different processing capacities.
@@
Figs. 13-14 show the job completion times for varying numbers of available servers and profiled processing capacities, respectively. These experiments are conducted with $\alpha=2$ and a system utilization of 75%, as we are particularly interested in the performance when resource contention is high. From Fig. 13, we observe that a larger number of available servers decreases job completion times. More available servers provide greater flexibility in task assignment, allowing tasks to be distributed in a manner that better balances the load among servers and reduces job completion times. Fig. 14 demonstrates that increasing processing capacity also reduces job completion times. This is expected, as higher processing capacities enable more tasks to run in parallel, thereby decreasing job completion times. Notably, the relative performance of the algorithms remains largely unchanged across different numbers of available servers and processing capacities.
@@ Acknowledgments
This work was supported in part by the Singapore Ministry of Education under Academic Research Fund Tier 2 Award MOE-T2EP20121-0005, the National Natural Science Foundation of China under Grants 62125206 and U20A20173, the National Key Research and Development Program of China under Grant 2022YFB4500100, and the Key Research Project of Zhejiang Province under Grant 2022C01145.